OpenStudy (anonymous):

What are the solutions for the equation? x ( to the second power) =11x - 24

OpenStudy (ash2326):

x^2=11x-24 x^2=11x-24 x^2-11x+24=0 find factors of 24 such that their sum is 3 and 8 are such that writing -11 as -3+ -8 x^2-3x-8x+24=0 x(x-3)-8(x-3)=0 (x-3)(x-8)=0 so x=3, 8